# Inhomogeneous mixing: From microscopic dynamics to mesoscopic staircases

### v1

@@ 2.3 Turbulent layer broadened by blobs/voids
Edge plasmas are usually strongly turbulent in linear plasma devices and toroidal fusion devices [44, 45, 46, 6]. Since blobs propagate outward and detach from the bulk plasma, while voids propagate inward and so stir the core plasma, it is natural to think of the roles of blobs/voids in the broadening of turbulent layer.
@@
Regarding the turbulent layer broadened by blobs, our recent experimental study shows that in high collisionality regime, the enhanced turbulence spreading induced by blobs can effectively broaden the power scrape-off width $\lambda_{q_{\parallel}}$ [47]. $\lambda_{q_{\parallel}}$ is one of the most important practical quantities of SOL since it controls the plasma power exhaust and thus governs the problems of heat removal and melting of plasma facing components (PFCs) [48]. The SOL collisionality parameter is used to distinguish regimes of SOL, including sheath-limited regime with low $\nu_{SOL,e}^{*}$ and conduction-limited regime with high $\nu_{SOL,e}^{*}$ [49]. Probe measurements of two discharges with different magnetic field $B_{t}$, plasma current $I_{p}$ and central line-averaged density $\overline{n}$ on J-TEXT are compared. In all these discharges, $\nu_{SOL,e}^{*}$ is 20-200, which is much larger than 10. Thus, the SOL is in the conduction-limited regime. For this regime, the power scrape-off width or power decay length $\lambda_{q_{\parallel}}\approx 2/7\lambda_{T}$. Here, $\lambda_{T}$ is the temperature decay length.
@@
It is constructive to compare the experimental widths $\lambda_{q,EXP}$ with the theoretical predictions from neoclassical heuristic drift model and turbulent broadening model. For neoclassical heuristic drift model, note that the preliminary HD (PHD) model is applied for zero-collisionality [50], while the generalized HD (GHD) model is applied for finite collisionality [51]. For turbulent broadening (TB) model, the impact of turbulence spreading on SOL broadening is considered [7, 52]. Turbulence intensity $E$ sets the power scrape-off layer width beyond the PHD width as $\lambda_{q,TB}=(\lambda_{q,PHD}^{2}+{\tau_{\parallel}}^{2}E)^{1/2}$. Here, $\tau_{\parallel}=L_{\parallel}/c_{s}$ is the SOL residence time of heat. As an estimate, the turbulence intensity is given by $E=V_{I}^{2}\tau_{ac}^{2}/\tau_{\parallel}^{2}$ [7]. Here, “mean jet velocity” $V_{I}$ represents the velocity of turbulence spreading [38], and $\tau_{ac}$ is the auto-correlation time of density fluctuations.
@@
Table 1 presents the experimental measured power scrape-off widths $\lambda_{q,EXP}$ and theoretical predicted widths for the two discharges with different blobs’ dynamics. The temporal and spatial features of blobs, including burst rate $\gamma_{b}$, lifetime $\tau_{life}$, radial propagation speed $v_{prop}$ and characteristic radial length $l_{r}$, are obtained via the routinely used conditional average methods [7, 19, 17]. Compared with the case of lower SOL collisionality $\nu_{SOL,e}^{*}\sim 33$ with weaker turbulence spreading, for the case of higher SOL collisionality $\nu_{SOL,e}^{*}\sim 92$ with stronger turbulence spreading, the blobs’ burst rate $\gamma_{b}$ increases by 1.5 times, lifetime $\tau_{life}$ increases by 1.4 times, radial propagation speed $v_{prop}$ increases by 9.0 times, characteristic radial length $l_{r}$ increases by 1.1 times, and the experimental measured power scrape-off widths $\lambda_{q,EXP}$ correspondingly increases by 1.7 times. Note that for the case of higher $\nu_{SOL,e}^{*}$ with more active blobs, the preliminary heuristic drift model ($\lambda_{q,PHD}\sim 0.194\ cm$) obviously underestimates power scrape-off width ($\lambda_{q,EXP}\sim 0.341\ cm$) and the generalized heuristic drift (GHD) model ($\lambda_{q,GHD}\sim 1.127\ cm$) significantly overestimates power scrape-off width, while the turbulence broadening model ($\lambda_{q,TB}\sim 0.350\ cm$) is in good agreement with the experimental measurement. The result above suggests the important roles of blobs and turbulence spreading in the broadening of power scrape-off width in the high collisionality regime. This is also consistent with previous experimental and simulation work [53, 54, 55].
Regarding the turbulent layer broadened by voids, a new theoretical model has demonstrated that the heretofore ignored process of voids emission can drive a broad turbulent layer of width $\sim 100\rho_{s}$ for typical parameters ($\rho_{s}$ is the ion sound speed gyroradius) [9]. Since the voids are closely related to inward turbulence spreading, this once again evokes the previous relevant suggestion that models of the pedestal structure or transport barrier should include the influence of coherent structures and turbulence spreading [56]. More experimental verification should be carried out in this aspect.
@@ 2.4 Interactions between blobs/voids and zonal flow
The interactions between blobs/voids dynamics and sheared flow are interesting. On the one hand, the $E\times B$ sheared flow can cause the decorrelation, distortion and breaking of blobs/voids and regulate the formation of blobs/voids and their transport [57, 58]. Recent experimental studies indicate that improvement of particle confinement and higher operational density are achieved in positive biasing discharges of J-TEXT as the outward motion of the blobs is blocked and the turbulent transport is reduced with stronger externally driven sheared flow [59, 60]. The turbulence spreading induced by blobs is found to decrease as the $E\times B$ shearing rate increases, while the turbulence spreading induced by voids is insensitive to $E\times B$ shearing rate, in the experiments which scan line-averaged density and plasma current approaching the Greenwald density limit [27].
@@
Figure 5: (a) Time series of the relative density fluctuations $\delta n/n$ (top) and of the perpendicular velocity (bottom); (b) wavelet power of $\delta n/n$ (black solid line) and of the poloidal-velocity fluctuations near the GAM frequency (green dashed line) in the plasma edge. Vertical red lines denote the times when voids are present. Reprinted from Sladkomedova et al. [6]. Copyright 2023 Cambridge University Press.
@@
On the other hand, radially propagating blobs/voids can excite drift waves and drive intermittent zonal flow [9, 61]. This result could qualitatively explain the strong correlation between zonal flow power bursts and the detection of voids measured by BES diagnostic in experiments on MAST spherical tokamak [6]. As shown in Figure 5(a), shortly after the voids of the relative density fluctuations $\delta n/n$ emerge, the perpendicular velocity $v_{perp}$ (approximate to poloidal velocity $v_{\theta}$) exhibits a similar intermittent bursty behavior. A sharp peak in the cross correlation and a near zero phase of the cross spectra at $\sim 8$ kHz between two poloidally separated $v_{perp}$ channels suggest the existence of the poloidally symmetric coherent flows. These are called geodesic acoustic modes (GAM), which is a branch of zonal flow with finite frequency [62]. Gas puff imaging (GPI) measurements on the HL-2A tokamak, which track the blob motions, suggest that the frequency of blob/void structures are more likely to be higher than $10$ kHz [63].
@@
In Figure 5(b), the black solid curve shows the time series of the wavelet power of $\delta n/n$ ($P_{\delta n/n}$) integrated over the frequency range of 10–500 kHz, the green dashed curve shows the time series of the wavelet power of the perpendicular velocity fluctuations near the GAM frequency integrated over 6–9 kHz ($P_{GAM}$), and the vertical red lines denote the times when voids are present. It is found that the burst of density voids occurred almost simultaneously with the transient rise of the broadband turbulent fluctuation power $P_{\delta n/n}$. It is also evident that a transient rise of GAM power $P_{GAM}$ occurs $\sim 1$ ms after the burst of density voids. Besides, recent experimental observations in the plasma edge of TJ-K stellarator also support this physical mechanism [64]. Blobs in the SOL are found to cause potential fluctuations in a zonal band inside the separatrix in a unidirectional causality by using probe measurements and convergent cross mapping methods. These results indicate that blobs and voids can play an important role in driving intermittent zonal flow.
3 A global structure by inhomogeneous mixing: the $E\times B$ staircase
3.1 Introduction to the $E\times B$ staircase
@@
The $E\times B$ staircase refers to a globally self-organized $E\times B$ zonal flow layers, separated by a mesoscopic scale. These sheared zonal flow layers act as weak or permeable transport barriers, resulting in a corrugated pressure profile resembling a staircase. The treads and riser of a corrugated profile correspond to the strong and weak mixing regions, respectively, representing an inhomogeneous mixing [65] of pressure. It was first identified in the full-f flux-driven electrostatic and adiabatic electron gyrokinetic simulations using GYSELA and XGC1 [66, 67]. Figure 6 shows the the normalized temperature gradient profile whose peaks align with the locations of strong flow shear layers. The $E\times B$ staircase can be most generally characterized by two spatial scales $\delta^{\mathrm{flow}}$ and $\Delta$, though they may not be clearly distinguishable in some cases [68]. $\delta^{\mathrm{flow}}$ indicates the radial extent of each zonal flow layer, and $\Delta$ is the distance between adjacent zonal flow layers or the radial extent of the mixing zone.
@@
Figure 6: The $E\times B$ staircase observed in the GYSELA simulation. Reprinted from Dif-Pradalier et al. [66] (Copyright 2010 AIP) and Dif-Pradalier et al. [69] (Copyright 2015 AIP).
@@
The $E\times B$ staircase is a pattern spontaneously emerging in the near-marginal regime [67]. The near-marginal regime means a state of plasmas that wander around the instability threshold due to transport processes. The turbulence transport in this regime is distinct from the expectations of commonly used reduced models based on the local diffusive paradigm [70, 71]. In particular, the Kubo number $Ku\equiv v\tau_{\mathrm{corr}}/\lambda_{\mathrm{corr}}=\tau_{\mathrm{corr}}/\tau_{\mathrm{turn}}>1$ is observed in a near-marginal regime [71]. The Kubo number is defined as the ratio between the distance traveled by a particle with typical velocity $v$ for the correlation time $\tau_{\mathrm{corr}}$ of an ambient potential eddy and the correlation length $\lambda_{\mathrm{corr}}$ of the eddy (or it can be written as the ratio between the correlation time $\tau_{\mathrm{corr}}$ and the eddy turnover time $\tau_{\mathrm{turn}}=\lambda_{\mathrm{corr}}/v$). When $Ku<1$, which is usual criterion for validity of a quasilinear theory, the transport is more like diffusion by random kicks by the rapidly decorrelating potential eddies. When $Ku>1$, the particles would have a sufficient time to interact with the potential eddies and their transport can exhibit other behaviors including either a sub-diffusion by being trapped within a localized eddy or a super-diffusion by traveling along a streamer [72] or the radially successively spreading [73] or phase matching [74] eddies. The latter can appear as avalanches, or ballistic propagation of blobs and voids whose size distribution follows the self-similar or scale-invariant power-law [40, 41, 42, 43], when a multitude of metastable states exists as in the self-organized criticality (SOC) system [70, 36, 75]. The $E\times B$ staircase forms as avalanches are regulated within $\Delta$ of spontaneously emerging zonal flow layers.
@@
The co-existence of avalanches and zonal flow layers makes the turbulence transport in the near-marginal regime more complicated. Understanding the $E\times B$ staircase would enable closing the gap between self-consistent – but expensive – gyrokinetic simulations and reduced – but practical – models [71]. The rest of this Section 3 presents a brief summary of rare experimental observations and characterizations of the $E\times B$ staircase, aiming to enhance its understanding and to encourage further experimental research. While the experimental findings may depend on a specific device configuration and plasma scenario (see Table 2), they can also reveal different aspects of the $E\times B$ staircase.
3.2 Observations of the $E\times B$ staircase
@@
In a broad sense, the $E\times B$ staircase is a set of weak transport barriers by localized multiple shear flow layers extending through the whole plasma. The conventional methods to identify a macroscopic transport barrier in experiments are based on one dimensional profiles of the measured pressure or the estimated radial electric field. However, identifying weak transport barriers only with one dimensional measurement of profiles may be challenging because the weak profile corrugation would not be clearly distinguished from measurement noises. Nonetheless, several approaches have been implemented in Tore Supra [69, 76], HL-2A [77], DIII-D [78], KSTAR [42, 79], TJ-II [80], and W7-X [81] plasmas, to overcome those limitations, as described below. The experimental conditions are summarized in Table 2. Results of different experiments may represent different aspects of the $E\times B$ staircase under the same physical origin, or it might be also possible that they are originated from completely different physical processes.
Table 2: Experiments on mesoscopic transport barriers
Device $R_{0}$ [m] $a/R$ $B_{0}$ [T] $I_{p}$ [MA] Heating [MW] Instability $\nu_{*}$ $\rho_{*}^{-1}$
@@
Figure 7: (a) The reduction of the radial correlation length near the shear flow extrema in the GYSELA simulation. (b) The measured correlation length profile, indicating multiple minima locations corresponding to the expected shear flow layers ($S_{0}$$S_{3}$) in the Tore Supra experiment. Reprinted from Dif-Pradalier et al. [69]. Copyright 2015 AIP. (c) The radial profiles of the temperature and density gradients, the radial correlation length, and the safety factor $q$ in the HL-2A experiment. Reprinted from Liu et al. [77]. Copyright 2021 AIP.
The first approach to identify the $E\times B$ staircase in experiments is to measure the variation of turbulence correlation length or the tilt of turbulent structures by the shear flow layers. For example, the local minima of the turbulence correlation length measurements along the radius were identified in Tore Supra L-mode plasmas [69, 76]. It was shown using the GYSELA data that the flow shear extrema are well correlated with the correlation length minima as shown in Figure 7(a). Due to a possible meandering behavior of the $E\times B$ staircase [69], the fast-sweeping X-mode reflectometry on Tore Supra was used to get an almost instantaneous radial measurements of turbulent fluctuations to obtain the correlation length profile. Figure 7(b) shows that multiple local minima ($S_{0}$$S_{3}$) were identified and they are not closely associated with the rational $q$ surfaces except $S_{0}$. Note that measurements were done in the MHD-free plasmas to avoid unwanted effects in the analyses. The reversed tilt angle of the time-radius correlation function was also provided as further evidence of the shear flow layers [76].
@@
The second approach is to improve the credibility of the measured radial corrugation in kinetic profiles using additional information from multi transport channels, fluctuations, or two-dimensional data. Figure 7(c) shows measurements of both electron temperature and density gradients, exhibiting multiple peaks aligned with the expected location of the shear flow layers from the turbulence analysis in the HL-2A plasmas [77]. A region of relatively enhanced mixing (more efficient turbulence spreading), which corresponds to a larger correlation length region in Figure [77](c), was identified in between the peaks of kinetic profiles (the transport barriers or suppressed mixing regions).
@@
In DIII-D H-mode edge plasmas [78], the axisymmetric two-step profiles were observed in multi channels such as density and electron and ion temperatures as shown in Figure 8(a). It was obtained by ensemble averaging over the grassy edge localized mode phases achieved by the $n=3$ resonant magnetic perturbation field. Repeated transport bifurcations between the two-step staircase and the normal single-step pedestal were observed.
@@
Figure 8: (a) The radial profiles of the electron and ion temperature, density, and $E\times B$ rotation frequency in the DIII-D experiment. Reprinted from Ashourvan et al. [78]. Copyright 2019 AIP. (b) The two-dimensional measurement of the radial corrugation of electron temperature in the KSTAR experiment. Reprinted from Choi et al. [42]. Copyright 2019 IOP.

### v2

@@ 2.3 Turbulent layer broadening by blobs/voids
Edge plasmas are usually strongly turbulent in linear plasma devices and toroidal fusion devices [45, 46, 47, 6]. Since blobs propagate outward and detach from the bulk plasma, while voids propagate inward and so stir the core plasma, it is natural to consider the roles of blobs/voids in the broadening of turbulent layer.
@@
Regarding the turbulent layer broadening by blobs, our recent experimental study shows that in high collisionality regime, the enhanced turbulence spreading induced by blobs can effectively broaden the power scrape-off width $\lambda_{q_{\parallel}}$ [48]. $\lambda_{q_{\parallel}}$ is one of the most important practical quantities of the SOL since it controls the plasma power exhaust and thus governs the problems of heat removal and melting of plasma facing components (PFCs) [49]. The SOL collisionality parameter is used to distinguish the regimes of the SOL, including the sheath-limited regime with low $\nu_{SOL,e}^{*}\approx 10^{-16}\frac{n_{LCFS}L_{\parallel}}{T_{\mathrm{e},LCFS}^{2}}$ and the conduction-limited regime with high $\nu_{SOL,e}^{*}$ [50]. Here, $n_{LCFS}$ is the plasma density at the last closed flux surface (LCFS), $T_{e,LCFS}$ is the electron temperature at the LCFS, and $L_{\parallel}$ is the connection length of magnetic field lines. Probe measurements of two discharges with different toroidal magnetic field $B_{\phi}$, plasma current $I_{p}$ and central line-averaged density $\overline{n}$ on J-TEXT are compared. In all these discharges, $\nu_{SOL,e}^{*}$ is 20-200, which is much larger than 10. Thus, the SOL is in the conduction-limited regime. For this regime, the power scrape-off width or power decay length $\lambda_{q_{\parallel}}\approx 2/7\lambda_{T}$. Here, $\lambda_{T}$ is the temperature decay length.
@@
It is constructive to compare the experimental widths $\lambda_{q,EXP}$ with the theoretical predictions from neoclassical heuristic drift model and turbulent broadening model. For neoclassical heuristic drift (HD) model, note that the preliminary HD (PHD) model is applied for zero collisionality [51], while the generalized HD (GHD) model is applied for finite collisionality [52]. For turbulent broadening (TB) model, the impact of turbulence spreading on SOL broadening is considered [7, 53]. Turbulence intensity $E$ sets the power scrape-off layer width beyond the PHD width as $\lambda_{q,TB}=(\lambda_{q,PHD}^{2}+{\tau_{\parallel}}^{2}E)^{1/2}$. Here, $\tau_{\parallel}=L_{\parallel}/c_{s}$ is the SOL residence time of heat where $c_{s}=(T_{\mathrm{e}}/m_{\mathrm{i}})^{1/2}$ is the ion sound speed. As an estimate, the turbulence intensity is given by $E=V_{I}^{2}\tau_{ac}^{2}/\tau_{\parallel}^{2}$ [7]. Here, “mean jet velocity” $V_{I}$ represents the velocity of turbulence spreading [38], and $\tau_{ac}$ is the auto-correlation time of density fluctuations.
@@
Table 1 presents the experimental measured power scrape-off widths $\lambda_{q,EXP}$ and theoretical predicted widths for the two discharges with different blobs’ dynamics. The temporal and spatial features of blobs, including burst rate $\gamma_{b}$, lifetime $\tau_{life}$, radial propagation speed $v_{prop}$ and characteristic radial length $l_{r}$, are obtained via the routine conditional average methods [7, 19, 17]. Compared with the case of lower SOL collisionality $\nu_{SOL,e}^{*}\sim 33$ with weaker turbulence spreading, in the case of higher SOL collisionality $\nu_{SOL,e}^{*}\sim 92$ with stronger turbulence spreading, the blobs’ burst rate $\gamma_{b}$ increases by 1.5 times, lifetime $\tau_{life}$ increases by 1.4 times, radial propagation speed $v_{prop}$ increases by 9.0 times, characteristic radial length $l_{r}$ increases by 1.1 times, and the experimental measured power scrape-off widths $\lambda_{q,EXP}$ correspondingly increases by 1.7 times. Note that for the case of higher $\nu_{SOL,e}^{*}$ with more active blobs, the preliminary heuristic drift model ($\lambda_{q,PHD}\sim 0.194\ cm$) obviously underestimates power scrape-off width ($\lambda_{q,EXP}\sim 0.341\ cm$) and the generalized heuristic drift (GHD) model ($\lambda_{q,GHD}\sim 1.127\ cm$) significantly overestimates power scrape-off width, while the turbulence broadening model ($\lambda_{q,TB}\sim 0.350\ cm$) is in good agreement with the experimental measurement. The result above suggests the important roles of blobs and turbulence spreading in the broadening of power scrape-off width in the high collisionality regime. This is also consistent with previous experimental and simulation work [54, 55, 56].
Recently, a new theoretical model has demonstrated that the heretofore ignored process of void emission can drive a broad turbulent layer of width $\sim 100\rho_{s}$ for typical parameters ($\rho_{s}$ is the ion sound speed gyroradius) [9]. Since the voids are closely related to inward turbulence spreading, this once again evokes the previous relevant suggestion that models of the pedestal structure or transport barrier should include the influence of coherent structures and turbulence spreading [57]. More experimental validation of this prediction should be carried out.
@@ 2.4 Interactions between blobs/voids and zonal flow
The interactions between blobs/voids and sheared flow are complex. On the one hand, the $\mathbf{E}\times\mathbf{B}$ sheared flow can cause the decorrelation, distortion, and breaking of blobs/voids, and regulate the formation of blobs/voids and their transport [58, 59]. Recent experimental studies indicate that improvement of particle confinement and higher operational density are achieved in positive biasing discharges of J-TEXT as the outward motion of the blobs is blocked, and the turbulent transport is reduced by stronger externally driven sheared flow [60, 61]. In the experiments which scan line-averaged density and plasma current approaching the Greenwald density limit, the turbulence spreading induced by blobs is found to decrease as the $\mathbf{E}\times\mathbf{B}$ shearing rate increases, while the turbulence spreading induced by voids is insensitive to $\mathbf{E}\times\mathbf{B}$ shearing rate [27].
@@
Figure 5: (a) Time series of the relative density fluctuations $\delta n/n$ (top) and of the perpendicular velocity (bottom); (b) wavelet power of $\delta n/n$ (black solid line) and of the poloidal velocity fluctuations near the GAM frequency (green dashed line) in the plasma edge. Vertical red lines denote the times when voids are present. Reprinted from Sladkomedova et al. [6]. Copyright 2023 Cambridge University Press.
@@
On the other hand, radially propagating blobs/voids can excite drift waves and drive intermittent zonal flow [9, 62]. This result could qualitatively explain the strong correlation between zonal flow power bursts and the detection of voids measured by BES diagnostic in experiments on MAST spherical tokamak [6]. As shown in Figure 5(a), shortly after the voids of the relative density fluctuations $\delta n/n$ emerge, the perpendicular velocity $v_{perp}$ (approximate to poloidal velocity $v_{\theta}$) exhibits a similar intermittent bursty behavior. A sharp peak in the cross correlation and a near zero phase of the cross spectra at $\sim 8$ kHz between two poloidally separated $v_{perp}$ channels suggest the existence of the poloidally symmetric coherent flows. These are called geodesic acoustic modes (GAMs), which are a branch of zonal flow with finite frequency [63]. Gas puff imaging (GPI) measurements on the HL-2A tokamak, which track the blob motions, suggest that the frequency of blob/void structures are more likely to be higher than $10$ kHz [64].
@@
In Figure 5(b), the black solid curve shows the time series of the wavelet power of $\delta n/n$ ($P_{\delta n/n}$) integrated over the frequency range of 10–500 kHz, the green dashed curve shows the time series of the wavelet power of the perpendicular velocity fluctuations near the GAM frequency integrated over 6–9 kHz ($P_{GAM}$), and the vertical red lines denote the times when voids are present. It is found that the burst of density voids occurred almost simultaneously with the transient rise of the broadband turbulent fluctuation power $P_{\delta n/n}$. It is also evident that a transient rise of GAM power $P_{GAM}$ occurs $\sim 1$ ms after the burst of density voids. Besides, recent experimental observations in the plasma edge of TJ-K stellarator also support this physical mechanism [65]. Blobs in the SOL are found to cause potential fluctuations in a zonal band inside the separatrix in a unidirectional causality by using probe measurements and convergent cross mapping methods. These results indicate that blobs and voids can play an important role in driving intermittent zonal flow. Moreover, formation of such void or blob driven zonal flows can initiate staircase formation, linking microphysics of blob/void and mesoscopic structure formation. It is to staircases and other mesoscopic phenomena that we turn in the following section.
3 Mesoscopic structure by inhomogeneous mixing: the $\mathbf{E}\times\mathbf{B}$ staircase
3.1 Introduction to the $\mathbf{E}\times\mathbf{B}$ staircase
@@
The $\mathbf{E}\times\mathbf{B}$ staircase refers to a self-organized $\mathbf{E}\times\mathbf{B}$ zonal flow layers, separated by a mesoscopic scale. These sheared zonal flow layers act as weak or permeable transport barriers, resulting in a corrugated pressure profile resembling a staircase. The treads and riser of a corrugated profile correspond to the strong and weak mixing regions, respectively, representing an inhomogeneous mixing [66] of pressure. It was first identified in the full-f flux-driven electrostatic and adiabatic electron gyrokinetic simulations using GYSELA and XGC1 [67, 68]. Figure 6 shows the normalized temperature gradient profile whose peaks align with the locations of strong flow shear layers. The $\mathbf{E}\times\mathbf{B}$ staircase can be most generally characterized by two spatial scales $\delta^{\mathrm{flow}}$ and $\Delta$. $\delta^{\mathrm{flow}}$ indicates the radial extent of each zonal flow layer, and $\Delta$ (tread width) is the distance between adjacent zonal flow layers or the radial extent of the mixing zone. These two may not be clearly distinguishable in some cases though [69].
@@
Figure 6: The $\mathbf{E}\times\mathbf{B}$ staircase observed in the GYSELA simulation with the ion collisionality $\nu_{*}=0.05$ and the normalized ion gyroradius $\rho_{*}=\rho_{\mathrm{i}}/a=1/256$. Reprinted from Dif-Pradalier et al. [67] (Copyright 2010 APS) and Dif-Pradalier et al. [70] (Copyright 2015 APS).
@@
The $\mathbf{E}\times\mathbf{B}$ staircase is a pattern which spontaneously emerges in the near-marginal regime [68]. The near-marginal regime means a state of the plasma which fluctuates close to the instability threshold. The near-marginal state is maintained by transport processes. The turbulence transport in this regime is distinct from the expectations of commonly used reduced models based on the local diffusive paradigm [71, 72]. In particular, large Kubo number $Ku>1$ is predicted in a near-marginal regime [72]. The Kubo number [73] can be written as the ratio of the correlation time $\tau_{\mathrm{corr}}$ to the eddy turnover time $\tau_{\mathrm{turn}}=\lambda_{\mathrm{corr}}/\tilde{v}$. When $Ku<1$, which is the usual criterion for validity of a quasilinear theory, the transport is more like diffusion by random kicks by the rapidly decorrelating potential eddies. When $Ku>1$, the particles would have sufficient time to interact with the potential eddies and their transport can exhibit other behaviors including either a sub-diffusion by being trapped within a localized eddy or a super-diffusion by traveling along a streamer [74], or the radially successively-spreading [36, 7] or phase-matching [75] eddies. The latter can appear as avalanches, or ballistic propagation of (density or temperature) blobs and voids whose size distribution follows the self-similar or scale-invariant power-law [40, 41, 42, 43], when a multitude of metastable states exists as in the self-organized criticality (SOC) system [71, 36, 76]. The $\mathbf{E}\times\mathbf{B}$ staircase forms as avalanches are regulated within $\Delta$ of spontaneously emerging zonal flow layers.
@@
The co-existence of avalanches and zonal flow layers makes the turbulence transport in the near-marginal regime more complicated. Understanding the $\mathbf{E}\times\mathbf{B}$ staircase would enable closing the gap between self-consistent – but expensive – gyrokinetic simulations and reduced – but practical – models [72]. The rest of this Section 3 presents a brief summary of rare experimental observations and characterizations of the $\mathbf{E}\times\mathbf{B}$ staircase. The aim is to enhance its understanding and to encourage further experimental research.
3.2 Observations of the $\mathbf{E}\times\mathbf{B}$ staircase
@@
In a broad sense, the $\mathbf{E}\times\mathbf{B}$ staircase is an array of weak transport barriers formed by localized multiple shear flow layers extending through the whole plasma. The conventional methods to identify a macroscopic transport barrier in experiments are based on one dimensional profiles of the measured pressure or the estimated radial electric field. However, identifying weak transport barriers only with one dimensional measurement of profiles may be challenging. The weak profile corrugation would not be clearly distinguished from measurement noise. To overcome this limitation, different approaches have been implemented in Tore Supra [70, 77], HL-2A [78], DIII-D [79], KSTAR [42, 80], TJ-II [81], and W7-X [82] plasmas as described below. The experimental conditions are summarized in Table 2. While the experimental findings may depend on a specific device configuration and plasma scenario, they might also reveal different aspects of the $\mathbf{E}\times\mathbf{B}$ staircase.
Table 2: Experiments on mesoscopic transport barriers. Key plasma parameters and experiment condition are listed: plasma major radius $R_{0}$, minor radius $a$, toroidal magnetic field strength (at $R_{0}$) $B_{\phi}$, plasma current $I_{p}$, heating methods (radio frequency (RF) resonance heating; neutral beam injection (NBI); electron cyclotron wave resonance heating (ECH)), dominant micro-instabilities (ion temperature gradient mode (ITG) with $\eta_{\mathrm{i}}\equiv L_{n}/L_{T_{\mathrm{i}}}$ drive where $L_{n}$ and $L_{T_{\mathrm{i}}}$ are the density and ion temperature gradient scale lengths, respectively; trapped electron mode (TEM); electron temperature gradient mode (ETG)), the ion collisionality $\nu_{*}$, and the inverse normalized ion gyroradius $\rho_{*}^{-1}=a/\rho_{\mathrm{i}}$.
Device $R_{0}$ [m] $a/R$ $B_{\phi}$ [T] $I_{p}$ [MA] Heating [MW] Instability $\nu_{*}$ $\rho_{*}^{-1}$
@@
Figure 7: (a) The reduction of the radial correlation length near the shear flow extrema in the GYSELA simulation. Black, red, and green lines represent the radial profile of the correlation length normalized by the ion gyroradius, $L_{c}/\rho_{0}$, at different normalized times $a/c_{s}$ where $a$ is the minor radius and $c_{s}=(T_{\mathrm{e}}/m_{\mathrm{i}})^{1/2}$ is the ion sound speed. (b) The measured correlation length profile, showing multiple minima locations expecting the shear flow layers ($S_{0}$$S_{3}$) in the Tore Supra experiment. Reprinted from Dif-Pradalier et al. [70]. Copyright 2015 APS. (c) The radial profiles of the temperature and density gradients, the radial correlation length, and the safety factor $q$ in the HL-2A experiment. Reprinted from Liu et al. [78]. Copyright 2021 AIP.
The first approach to identify the $\mathbf{E}\times\mathbf{B}$ staircase in experiments is to measure the variation of turbulence correlation length or the tilt of turbulent structures by the shear flow layers. For example, the local minima of the turbulence correlation length measurements along the radius were identified in Tore Supra L-mode plasmas [70, 77]. It was shown using the GYSELA data that the flow shear extrema are well correlated with the correlation length minima as shown in Figure 7(a). Due to a possible meandering behavior of the $\mathbf{E}\times\mathbf{B}$ staircase [70], the fast-sweeping X-mode reflectometry on Tore Supra was used to get an almost instantaneous radial measurements of turbulent fluctuations to obtain the correlation length profile. Figure 7(b) shows that multiple local minima ($S_{0}$$S_{3}$) were identified, and they are not closely associated with the rational $q$ surfaces except $S_{0}$. Note that measurements were done in plasmas without MHD activity to avoid unwanted effects in the analyses. The reversed tilt angle of the time-radius correlation function was also provided as further evidence of the shear flow layers [77].
@@
The second approach is to improve the credibility of the measured radial corrugation in plasma profiles using additional information from multi transport channels, fluctuations, or two-dimensional data. Figure 7(c) shows measurements of both electron temperature and density gradients, exhibiting multiple peaks aligned with the expected location of the shear flow layers from the turbulence analysis in the HL-2A plasmas [78]. A region of relatively enhanced mixing (more efficient turbulence spreading), which corresponds to a larger correlation length region in Figure 7(c), was identified in between the peaks (the transport barriers or suppressed mixing regions) of plasma profiles.
@@
In DIII-D H-mode edge plasmas [79], the axisymmetric two-step profiles were observed in multi channels such as density and electron and ion temperatures as shown in Figure 8(a). It was obtained by ensemble averaging over the weak edge-localized-mode phases achieved by the resonant magnetic perturbation field. Repeated transport bifurcations between the two-step staircase and the normal single-step pedestal were observed.
@@
Figure 8: (a) The radial profiles of the electron temperature $T_{\mathrm{e}}$ and ion temperature $T_{\mathrm{i}}$, density $n_{\mathrm{e}}$, and $\mathbf{E}\times\mathbf{B}$ rotation frequency in the DIII-D experiment. Reprinted from Ashourvan et al. [79]. Copyright 2019 APS. (b) The two-dimensional (2D) measurement of the radial corrugation of normalized electron temperature variation $\frac{\delta T_{\mathrm{e}}}{\langle T_{\mathrm{e}}\rangle}=\frac{T_{\mathrm{e}}-\langle T_{\mathrm{e}}\rangle}{\langle T_{\mathrm{e}}\rangle}$ in the KSTAR experiment. Here, $\langle\rangle$ means the time average. Peaks (p1–p3) and valleys (v1–v2) appear as concentric jet-like patterns in 2D. $\Delta$ ($\Delta_{z}$) indicates the tread width of temperature staircase along the radial (vertical) direction. Reprinted from Choi et al. [42]. Copyright 2019 IOP.